OpenStudy (anonymous):

Convert the following equation from rectangular to polar form y = -3

OpenStudy (anonymous):

please help me! :)

OpenStudy (anonymous):

not really sure about it, but: \[\sqrt{x ^{2}+y ^{2}}=r\] \[arctg y/x=\alpha \] this are the change of coordinates equations. Just put -3 where y

myininaya (myininaya):

 \[\cos(\theta)=\frac{x}{r}\] \[\sin(\theta)=\frac{y}{r}\] => \[ r \cos(\theta)=x\] \[ r \sin(\theta)=y\] One of these equations tells you what to replace y with